OpenStudy (rebeccapink6398):

someone pls explain this question to me? i will medal and fan thanks so much for helping

OpenStudy (rebeccapink6398):

sorry for the split screen lol i was watching a show

OpenStudy (anonymous):

\[\sqrt[3]{125}=\left( 125 \right)^{\frac{ 1 }{ 3 }}=\left\{ \left( 5 \right)^3 \right\}^{\frac{ 1 }{ 3 }}\] can you solve it further?